how do you solve this equation into slope intercept form:
-2y-10+2x=0?
3x+20=-4y?
please help me!!

Answer by stanbon(57332)

how do you solve this equation into slope intercept form:
-2y-10+2x=0?
3x+20=-4y?
------------------
1st: -2(y-5-x) = 0
y-x-5 = 0
y = x+5
---------------------
2nd: 3x+20 = -4y
y = (-3/4)x - 5
============================
If this is a system of 2 equation:
Substitute to solve for "x":
x+5 = (-3/4)x-5
(7/4)x = -10
x = (4/7)*10
x = 40/7
-----------------------------
Solve for "y":
y = x + 5
y = (40/7) +(35/7)
y = 75/7
=============================
Cheers,
Stan H.